Suppose there are 4 defective items in a lot (collection) of 60 items. A sample of size 15 is taken at random and without repConstruct the probability table and graph the probability mass function.In the sample, at least one defective item is detected. What is the probability that at most two defective items are found?please show work..
Answer #1

Given X denote the number of defective items in the sample with probability p = 4/60 = 0.06667 and the sample size is 15.

So the distribution X ~ B(15,0.06667)

See the image for the probability table(left) and probability mass function(right).

III Binomial Distribution desmos Create Account ОГ Sign In << 11 0 + -0.6 4 Х B(n.p,x) 0 0.35526436 1 0.38064039 2 0.1903202

Now in last part, it is asked P(X<=2 | X > 1)

which is P(X<= 2  \bigcap X>1 ) / P(X>1) ==> P(X=2) /1 - P(X<=1)

P(X=2) = 0.190 and P(X<=1) = 0.355+0.380 = 0.735 (Note: The values are taken from probability table in image)

==> P(X=2) /1 - P(X<=1) = 0.190/ 1-0.735 =  0.7169